Find the time required for an investment of 5000 dollars to grow to 6000 dollars at an interest rate of 75 percent per year, compounded quarterly Round your answer to two decimal places Your answer istl Preview years Get help: Video Submit
Answer #1

Present Value = $5,000
Future Value = $6,000

Annual Interest Rate = 7.50%
Quarterly Interest Rate = 1.875%

Let time period be t quarters

Present Value * (1 + Quarterly Interest Rate)^Period = Future Value
$5,000 * 1.01875^t = $6,000
1.01875^t = 1.20
t * ln(1.01875) = ln(1.20)
t = 9.815

Number of quarters = 9.815
Number of years = 9.815 / 4
Number of years = 2.45